Q: Given the demand equation, Q = 12,000 - 10P2.
a. For this equation, write the expression for the point price elasticity of demand as a function of P
b. over what range of prices is the demand inelastic ?
1
Expert's answer
2018-02-15T13:56:08-0500
a) EpD = -10*P2/Q = -10 * P2*/(1200-10*P2)= P2/(P2-120)
b) Demand is inelastic when price is (0;60).